OpenStudy (anonymous):

evaluate the expression . (1/3 ) ^3 *( 4 +2 ) ^3 -( -6 ) ^3 ?

OpenStudy (anonymous):

(0.33)^3*(6)^3-(-6)^3 0.035937*216+216 7.762392+216=223.762392

OpenStudy (anonymous):

1/27*216+216=28/27*216=224

OpenStudy (anonymous):

charlie1092 i think you did a bit complicated things

OpenStudy (anonymous):

yea i put to many number i like your way better